OpenStudy (aaronandyson):

Prove that: sin^8 x - cos^8x = (sin^2(x) - cos^2(x))(1- 2sin^2(x)cos^2(x))

OpenStudy (aaronandyson):

@baru

OpenStudy (aaronandyson):

do I take the square of sin and cos then cube the whole bracket?

OpenStudy (aaronandyson):

\[[(\sin^2(x) - \cos^2(x)]^3\]

OpenStudy (baru):

no, do you know this identity\[a^2-b^2=(a+b)(a-b)\] ?

OpenStudy (baru):

lets work on the left side of the equaiton \[\sin^8(x) - \cos^8(x)\\=(\sin^4(x))^2-(\cos^4(x))^2\] ok?

OpenStudy (aaronandyson):

Okay.?

OpenStudy (baru):

\[(\sin^4(x))^2-(\cos^4(x))^2\\=(\sin^4(x)+\cos^4(x))\times(\sin^4(x)-\cos^4(x))\] ok? ( i used the identity i mentioned in my first post)

OpenStudy (aaronandyson):

okay.

OpenStudy (baru):

i will use the same identity again and simplify the second term \[(\sin^4(x)+\cos^4(x))\times \color{red}{(\sin^4(x)-\cos^4(x))}\\(\sin^4(x)+\cos^4(x))\times \color{red}{((\sin^2(x))^2-(\cos^2(x))^2)}\\(\sin^4(x)+\cos^4(x))\times \color{red}{(\sin^2(x)+\cos^2(x))(\sin^2(x)-\cos^2(x))}\]

OpenStudy (baru):

followed?

OpenStudy (aaronandyson):

yes..

OpenStudy (aaronandyson):

sin^2 x + cos^2 x is 1..?

OpenStudy (baru):

we know the identity \[\sin^1(x)+\cos^2(x)=1\] so the above expression becomes \[(\sin^4(x)+\cos^4(x)) \times \color{red}{1 \times(\sin^2(x)-\cos^2(x))}\]

OpenStudy (baru):

yes :)

OpenStudy (aaronandyson):

what next?

OpenStudy (aaronandyson):

do I square root the equation?

OpenStudy (baru):

no, now is the tricky part do you know the property \[(a-b)^2=a^2+b^2-2ab\] ?

OpenStudy (aaronandyson):

Yes.

OpenStudy (baru):

we have \[(\sin^4(x)+\cos^4(x)) \times(\sin^2(x)-\cos^2(x))\] add and subtract as shown \[(\sin^4(x)+\cos^4(x) \color{red}{-2\sin^2(x)\cos^2(x)+2\sin^2(x)\cos^2(x)}) \times(\sin^2(x)-\cos^2(x))\] ok?

OpenStudy (aaronandyson):

I didn't get that. :(

OpenStudy (baru):

we have this \[(\sin^4(x)+\cos^4(x)) \times(\sin^2(x)-\cos^2(x))\] look at the term i mark in red,\[(\color{red}{\sin^4(x)+\cos^4(x)}) \times(\sin^2(x)-\cos^2(x))\] i will add and subtract \(\color{blue}{2sin^2(x)cos^2(x)}\) \[(\color{red}{\sin^4(x)+\cos^4(x)} \color{blue}{-2\sin^2(x)\cos^2(x)+2\sin^2(x)\cos^2(x)}) \times(\sin^2(x)-\cos^2(x))\] ok?

OpenStudy (aaronandyson):

How'd you get 2sin^2(x)cos^2(x)?

OpenStudy (baru):

if you add and subtract the same number to an equation, it remains unchanged. ex: 2x + 3y = 8 add and subtract 11 on the left 2x + 3y +11 - 11 =8 the equation is unchanged because +11-11 =0

OpenStudy (baru):

if you are wondering why specifically 2cos^2sin^2, you will see the reason in the following steps

OpenStudy (baru):

ok?

OpenStudy (aaronandyson):

Okay?

OpenStudy (baru):

\[(\color{red}{\ \color{green}{[[}\sin^4(x)+\cos^4(x)} \color{blue}{+2\sin^2(x)\cos^2(x)\color{green}{]]}-2\sin^2(x)\cos^2(x)}) \times(\sin^2(x)-\cos^2(x))\] look at the terms i put inside the green brackets [[ ]] they are in the form \[a^2+b^2+2ab\] where \[a=\sin^2(x)\\b=\cos^2(x)\] so we can write it as \[(\color{red}{\ \color{green}{[[}\sin^2(x)+\cos^2(x)} \color{green}{]]^2} \color{blue}{-2\sin^2(x)\cos^2(x)}) \times(\sin^2(x)-\cos^2(x))\]

OpenStudy (baru):

ok?

OpenStudy (aaronandyson):

uh,okay?

OpenStudy (baru):

we know sin^2+cos^2=1 so this simplifies to \[(\color{red}{1} \color{blue}{-2\sin^2(x)\cos^2(x)}) \times(\sin^2(x)-\cos^2(x))\] which is the expression on the right side so LHS = RHS